OpenStudy (anonymous):

x^2+5x=0 solve by completing the square

OpenStudy (jtvatsim):

First write x^2 + 5x = 0. We focus on the left side first. To get x^2 with FOIL, I need an x in both parentheses. So, (x )(x ) To figure out what we need next, take the 5x term, and look at the 5. Then divide it by two, giving us 5/2. So, (x 5/2)(x 5/2) We have positive numbers so we will have to add. Then, (x + 5/2)(x + 5/2). When we FOIL we see, x^2 +5x + 25/4. Our original equation was x^2 + 5x = 0. So add 25/4 to both sides and get. x^2 + 5x + 25/4 = 25/4. But, we know that x^2 + 5x + 25/4 is just (x + 5/2)(x + 5/2) or (x + 5/2)^2. So, (x + 5/2)^2 = 25/4. Take the square root of both sides. x + 5/2 = plus or minus 5/2. So, x = 0 or x = -5.

OpenStudy (anonymous):

thanku